The inverse point of 1+i with respect to the circle |z-1|=2 is


1
Expert's answer
2021-05-11T15:15:34-0400

Formula for symmetric point with respect to the circle-

"|z-z_o|=r"


So, "z'=\dfrac{r^2}{\bar{z}-z_o}+z_o~~~~~~~~~-(1)"


Given circle is-


"|z-1|=2"  


"\text{ So, } z_o=1,r=2 \text{ and point }z=1+i"


Subsititue above values in (1)-


"z'=\dfrac{2^2}{\overline{1+i}-1}=\dfrac{4}{1-i-1}=\dfrac{4}{-i}=4i"


Hence The inverse is 4i.